# Walk Wisely on Graph: Knowledge Graph Reasoning with Dual Agents via Efficient Guidance-Exploration

### v3

@@ Introduction
Consistent with (Zhang et al. 2022), we employ TransE (Bordes et al. 2013) due to its efficiency in encoding the structural proximity information of KG to generate pre-trained entity embeddings, then we divide the original KG into $N$ clusters by utilizing K-means. In order for GIANT can easily walk on the cluster-level graph $\mathcal{G}^{c}$ while preserving the relation information of the original KG $\mathcal{G}$, we add a link to two clusters if there is at least one entity-level edge between them, which is detailed in Appendix E [^1] Our extended version is available at
https://arxiv.org/pdf/2408.01880
@@
where $\mathbf{W}_{1}^{c}\in\mathbb{R}^{4d\times 4d}$, $\mathbf{W}_{2}^{c}\in\mathbb{R}^{2d\times 4d}$, $\mathbf{W}_{1}^{e}\in\mathbb{R}^{6d\times 6d}$, $\mathbf{W}_{2}^{e}\in\mathbb{R}^{2d\times 6d}$ are the matrices of learnable weights to maintain dimension of history embedding. While $\mathbf{A}_{t}^{c}\in\mathbb{R}^{|A_{t}^{c}|\times 2d},\mathbf{A}_{t}^{e}\in\mathbb{R}^{|A_{t}^{e}|\times 2d}$ represent the embeddings of all next possible actions for GIANT and DWARF respectively.
@@
Constraint Reward.  Considering the constraint, that is, receiving high-quality guidance from GIANT, we introduce a metric for the state similarity between GIANT and DWARF, denoted as $\text{Sim}(s_{t}^{c},s_{t}^{e})$. It is calculated as the cosine similarity between the pre-trained embeddings of the current cluster and the current entity:
@@
the Lagrange multiplier is updated to make the guidance-exploration trade-off. Figure 3 gives an illustration: When the state similarity metric between GIANT and DWARF $D(s_{t}^{c},s_{t}^{e})$ is less than the threshold, the parameter $\theta_{\lambda}$ are optimized so that $\lambda(s_{t}^{e})$ increases to encourage GIANT to provide more guidance for DWARF. Conversely, the parameter $\theta_{\lambda}$ are optimized so that $\lambda(s_{t}^{e})$ decreases to encourage DWARF to explore in the constrictive space when the state similarity metric between GIANT and DWARF $D(s_{t}^{c},s_{t}^{e})$ is greater than the threshold. In practice, we utilize a cross-entropy loss to optimize $\theta_{\lambda}$:
@@
FULORA + TransE (Bordes et al. 2013) 72.5/49.1/33.4 69.4/45.6/25.4 89.2/57.2/50.7
@@ A.1  Baseline Methods
We compare the performance of FULORA with the following baselines [^2] Code to reproduce the main experiments is available at: https://github.com/KotoHanon/LOCOCO., including embedding-based, GNN-based and multi-hop based KG reasoning methods:
@@ Appendix D E  Proof of Dynamic Path Feedback
where $\hat{Q}^{*}(s^{c}_{t},a^{c}_{t})=Q^{*}(s_{t}^{c},a_{t}^{c})-\alpha D(s^{c}_{t},s^{c}_{\rm{target}})$. Equation 14 is the Bellman optimal equation in dynamic path feedback. It constructs a new Markov Decision Process (MDP), which we use $M^{\prime}$ to denote for distinction and $M$ for the original MDP. Thus, the optimal policy in $M^{\prime}$ is

### v4

@@ Introduction
Consistent with (Zhang et al. 2022), we employ TransE (Bordes et al. 2013) due to its efficiency in encoding the structural proximity information of KG to generate pre-trained entity embeddings, then we divide the original KG into $N$ clusters by utilizing K-means. In order for GIANT can easily walk on the cluster-level graph $\mathcal{G}^{c}$ while preserving the relation information of the original KG $\mathcal{G}$, we add a link to two clusters if there is at least one entity-level edge between them, which is detailed in Appendix E.
@@
where $\mathbf{W}_{1}^{c}\in\mathbb{R}^{4d\times 4d}$, $\mathbf{W}_{2}^{c}\in\mathbb{R}^{2d\times 4d}$, $\mathbf{W}_{1}^{e}\in\mathbb{R}^{4d\times 4d}$, $\mathbf{W}_{2}^{e}\in\mathbb{R}^{2d\times 4d}$ are the matrices of learnable weights to maintain dimension of history embedding. While $\mathbf{A}_{t}^{c}\in\mathbb{R}^{|A_{t}^{c}|\times 2d},\mathbf{A}_{t}^{e}\in\mathbb{R}^{|A_{t}^{e}|\times 2d}$ represent the embeddings of all next possible actions for GIANT and DWARF respectively.
@@
Constraint Reward.  Considering the constraint, that is, receiving high-quality guidance from GIANT, we introduce a metric for the state similarity between GIANT and DWARF, denoted as $\text{Sim}(s_{t}^{c},s_{t}^{e})$. It is calculated as the cosine similarity between the pre-trained embeddings of the current cluster [^1] Each cluster embedding is obtained by averaging all entity embeddings within it. A slight abuse of notation indicates that the $\mathbf{c}_{t}$ here is not the $\mathbf{c}_{t}$ mentioned in the policy network. In the concrete implementation, we treat the $\mathbf{c}_{t}$ in the policy network as a self-cascade of the $\mathbf{c}_{t}$ here. and the current entity:
@@
the Lagrange multiplier is updated to make the guidance-exploration trade-off. Figure 3 gives an illustration: When the state similarity metric between GIANT and DWARF $\text{Sim}(s_{t}^{c},s_{t}^{e})$ is less than the threshold, the parameter $\theta_{\lambda}$ are optimized so that $\lambda(s_{t}^{e})$ increases to encourage GIANT to provide more guidance for DWARF. Conversely, the parameter $\theta_{\lambda}$ are optimized so that $\lambda(s_{t}^{e})$ decreases to encourage DWARF to explore in the constrictive space when the state similarity metric between GIANT and DWARF $\text{Sim}(s_{t}^{c},s_{t}^{e})$ is greater than the threshold. In practice, we utilize a cross-entropy loss to optimize $\theta_{\lambda}$:
@@
FULORA + TransE (Bordes et al. 2013) 72.5/49.1/36.4 69.4/45.6/27.1 89.2/57.2/51.3
@@ A.1  Baseline Methods
We compare the performance of FULORA with the following baselines, including embedding-based, GNN-based and multi-hop based KG reasoning methods:
@@ A.2  Data Statistics
A.3  Implementation Details
We use the following software versions:
  • Ubuntu 24.04 LTS
  • Python 3.10
  • Pytorch 2.0.1
We conduct all experiments with a single NVIDIA GeForce 3080Ti GPU. Our experiments of all of 12 baselines are conducted on their official implementation provided by their respective authors. To reproduce the results of our model in Table 2 and Table 3, we report the empirically optimal crucial hyperparameters as shown in Table 4. On all datasets, the quantities of path rollouts in training and testing are 20 and 100, separatively. The core codes of FULORA are available at: https://github.com/KotoHanon/LOCOCO.
@@ Appendix D E  Proof of Dynamic Path Feedback
where $\hat{Q}^{*}(s^{c}_{t},a^{c}_{t})=Q^{*}(s_{t}^{c},a_{t}^{c})-\alpha\text{Sim}(s^{c}_{t},s^{c}_{\rm{target}})$. Equation 14 is the Bellman optimal equation in dynamic path feedback. It constructs a new Markov Decision Process (MDP), which we use $M^{\prime}$ to denote for distinction and $M$ for the original MDP. Thus, the optimal policy in $M^{\prime}$ is